OpenStudy (3214evergaming):

The total cost of a pair of jeans and three shirts is $114. The shirts each cost the same amount, and the jeans are $45. What is the price per shirt? $53 $23 $69 $159

OpenStudy (mathstudent55):

Let j = price of jeans and let s = price of shirt j + s + s + s = 114 j + 3s = 114 We are told that j = 45, so substitute 45 for j: 45 + 3s = 114 Now solve the equation for s. First, subtract 45 from both sides. Then divide both sides by 3.